6. What is the asymptotic solution to the recurrence relation T(n) = 3T(n/2)+n3 log(n)?

please explain
Answer #1
Master theorem:
-----------------
T(n) = aT(n/b) + f(n)
If f(n) = Θ(n^c) where c < Logb(a) then T(n) = Θ(n^Logb(a))
If f(n) = Θ(n^c) where c = Logb(a) then T(n) = Θ((n^c)(Log(n)))
If f(n) = Θ(n^c) where c > Logb(a) then T(n) = Θ(f(n))

Given
T(n) = 3T(n/2)+n3log(n)
T(n) = 3T(n/2)+Θ(n3log(n))
Where a = 3, b = 2, f(n) = Θ(n3log(n)) then c is greater than 3
Log2(3)
= 1.5849625007 < c

So, It satisfies rule 3

T(n) = Θ(f(n)) = Θ(n3log(n))